suppose that when the price increases from 7 to 14, the quantity supplied increases from 35 to 82. what is the elasticity of supply
Answer #1

Price elasticity of supply=% change in Quantity supplied/% change in Price

% change in Quantity supplied=(82-35/35)*100=134.2

% change in price=(14-7/7)*100=100

Price elasticity of supply=134.2/100

Price elasticity of supply=1.34

Price elasticity of supply is 1.34, which means supply is elastic